OpenStudy (anonymous):

Which line shows the first error in the solution?4x+8=2x- 10 (1) 4x – 2x +8 = -10 – 8 (2) 2x + 8= -10 (3) 2x + 8 – 8 = -10 -8 (4) 6x=-2 (5) 6x/6 = -2/6 (6) X= - 1/3 A. Line 2 B. Line 3 C. Line 4 D. Line 5

OpenStudy (anonymous):

HELLP

OpenStudy (anonymous):

eh.. so much work :/ cannt u do it urself and ill check ur answer?

OpenStudy (anonymous):

i tried already

OpenStudy (anonymous):

well first i dont understand the question lol can u eladorate?

OpenStudy (anonymous):

i need to know where the first mistake is at

OpenStudy (anonymous):

line 1. Because 4x + 8 = 2x - 10 (subtract 2x from each side) 4x - 2x + 8 = -10

OpenStudy (anonymous):

how is that a mistake..

OpenStudy (anonymous):

great i got that right thank you

OpenStudy (anonymous):

if you say so! u have a higher score than me anyway xD

OpenStudy (anonymous):

4x - 2x + 8 = -10 (this is the next step)....you do not subtract 8 till the next step.

OpenStudy (anonymous):

ah, ok i c now lol i didnt understand the question :s